OpenStudy (anonymous):

A rectangular prism has length thirteen meters, width six meters, and volume three hundred ninety cubic meters. What is the height of the prism?

OpenStudy (zpupster):

I will help you set up the equation the Volume of rectangular prism is V= l x w x h where v is volume l is length and w is width and h is height so lets punch in our numbers : 360 m ^3 = 13m x 6m x h can you solve it now????

OpenStudy (anonymous):

78?

OpenStudy (zpupster):

first lets multiply the right hand side 13 x 6=?

OpenStudy (anonymous):

78

OpenStudy (zpupster):

78 m ^2 ok

OpenStudy (zpupster):

now we have to isolate the h on one side can you do that??

OpenStudy (anonymous):

idk how i am new to this

OpenStudy (zpupster):

ok in isolating we will divide both sides by 78m^2

OpenStudy (zpupster):

so 360m^3/78m^2 = 78m^2(h)/ 78M^2

OpenStudy (zpupster):

the 78 cancel on the right leaving with just h and on the left 360/78 =??m

OpenStudy (anonymous):

4.615

OpenStudy (zpupster):

yes remember the units : 4.615 meters Good Job!!!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i need help with one more thing which is graphing the equality

OpenStudy (anonymous):

Graph the inequality three times n minus four is less than or equal to twenty-three.

OpenStudy (zpupster):

post that as a separate problem ice